# A Near Quantum Limited Sub-GHz TiN Kinetic Inductance Traveling Wave Parametric Amplifier Operating in a Frequency Translating Mode

### v1

@@ Abstract
We present the design and experimental characterization of a kinetic-inductance traveling-wave parametric amplifier (KI-TWPA) for sub-GHz frequencies. KI-TWPAs amplify signals through nonlinear mixing processes supported by the nonlinear kinetic inductance of a superconducting transmission line. The device described here utilizes a compactly meandered TiN microstrip transmission line to achieve the length needed to amplify sub-GHz signals. It is operated in a frequency translating mode where the amplified signal tone is terminated at the output of the amplifier, and the idler tone at approximately 2.5 GHz is brought out of the cryostat. By varying the pump frequency, a gain of up to 22 dB was achieved in a tunable range from about 450 to 850 MHz. Use of TiN as the nonlinear element allows for a reduction of the required pump power by roughly an order of magnitude relative to NbTiN, which has been used for previous KI-TWPA implementations. This amplifier has the potential to enable high-sensitivity and high-speed measurements in a wide range of applications, such as quantum computing, astrophysics, and dark matter detection.
@@ I Introduction
Kinetic inductance traveling-wave parametric amplifiers (KI-TWPAs)[1] with near quantum-limited noise performance and wideband gain have been developed recently and used for the readout of low-temperature detectors and qubits [2, 3, 4, 5]. The application of KI-TWPAs to direct dark matter searches is also gaining attention [6, 7]. The majority of these amplifiers have been designed for the 4-8 GHz band. However, a few technologies and experiments in quantum information science, astronomy, and fundamental physics can benefit from a near-quantum-noise-limited amplifier with gain at sub-GHz frequencies. For example, single-shot readout of spin-qubits operating below 1 GHz may be possible using a quantum-limited sub-GHz amplifier [8, 9, 10, 11]. Reducing the readout noise of nano-calorimeters and nano bolometers would allow them to be sensitive to single photons in the microwave regime [12, 13, 14, 15, 16, 17]. Circuit quantum cavity electrodynamics with micro-mechanical hybrid systems can also benefit from sub-GHz quantum limited amplifiers [18, 19]. Far-IR MKID detectors such as the prototypes for PRIMA far-IR PRobe [20] and future mid-IR MIKIDs operating in the single photon counting regime would require low noise readout amplifiers operating around 1 GHz [21].
@@
This article presents results on a wide bandwidth sub-GHz KI-TWPA with sensitivity close to the quantum limit. The design of the amplifier draws on our earlier work[28, 29], but introduces two new innovations. The first departure from previous work is the use of a TiN microstrip line in place of NbTiN as the nonlinear transmission line, which significantly reduces the required pump power. Reduced pump power facilitates thermal control of the experimental setup at the low temperatures required to avoid thermal noise. For example, at 500 MHz, 1 photon corresponds to $hf/k_{\mathrm{B}}=24$ mK. KI-TWPA designs based on 4-wave mixing (4WM) typically needed a pump power of $\sim 10\,\mu$W. While the KI-TWPA itself only dissipates only a fraction of the pump power, it is harder to control dissipation in the surrounding circuitry, including microwave isolator. The TiN device described here requires only $\sim 0.3\,\mu$W, which causes minimal heating at the base temperature of a dilution refrigerator.
@@ II Device Design
Similar to KI-TWPAs produced previously in our laboratory, the device reported here is based on a nonlinear transmission line (NTRL) structure with an inverted microstrip line geometry [28, 30, 29]. The microstrip conductor layer is patterned from a 50 nm thick TiN film. The dielectric of the microstrip is a 100 nm thick amorphous Si layer, and the sky plane is 200 nm thick niobium. The width of the microstrip line is 320 nm. Every 15 $\mu$m  a pair of microstrip open stubs of the same width and average length 120 $\mu$m  are connected to the central microstrip line. The NTRL structure is tightly meandered, as shown in fig. 1a., across a 25 mm long chip. The 45 cm total length of the NTRL is several times longer than KI-TWPAs designed for > 4 GHz in order to roughly maintain the number of wavelengths for the sub-GHz signal frequencies.
@@
Figure 2: a) Plot of the gain of the device with an applied -35.8 dBm pump tone at 1.6 GHz, measured using a VNA. The input and output frequencies are the same for this measurement. The dashed line is the gain calculated by integrating coupled mode equations.[28, 31] b) The device may also be operated as a frequency translating amplifier using a set of diplexers. At its input, the signal band is admitted while the device sees a termination over the idler band. At the output, the idler band is transmitted, and the signal band is terminated.
@@ III Performance
Figure 3: Measurements in subfigures a,b, and c were taken at 200 mK with a 20 dB attenuator at the 4 K stage. A 20 dB directional coupler was used at room temperature to couple the pump and signal tones into the cryostat. a) Measured signal band gain for a range of pump frequencies. b) Corresponding idler band gain. c) Plot of the measured signal gain at 200 mK as a function of pump power for a pump tone at 1.6 GHz. The maximum gain near 20 dB corresponds to a pump power of 330 nW. d) Plot of measured output power at the spectrum analyzer as a function of signal tone power for two pump tone powers at a pump frequency of 1.5 GHz. The measurements were taken in a dilution refrigerator at 10 mK in the frequency translating gain setup of Fig. 5.
Applying a pumping tone below the stop band frequency results in parametric amplification via 4WM for weaker signal tones. The typical gain shown in fig. 2.a has three disjointed bands where gain is produced, similar to the device described in [29], but with a wider separation between bands. The bands reflect the frequencies at which the phase-matching criterion for 4WM is approximately satisfied:
@@
where the $k(\omega_{s,i,p})$ are the propagation constants at the signal, idler and pump frequencies, and $2\omega_{p}=\omega_{s}+\omega_{i}$. The dashed line in the figure shows the gain calculated by integrating the coupled mode equations [28, 31] including third harmonic and upper sideband generation (tones at $3\omega_{p}$, $2\omega_{p}+\omega_{s}$ and $2\omega_{p}+\omega_{i}$. The film penetration depth in the model was adjusted to 625 nm to match the measured stop band frequency. To match the overall gain, the pump current was set to $I_{p}/I_{\ast}=0.117$, corresponding to 300 nW, which is close to the estimate of the experimental pump power.
@@
Fig. 3.c shows the response of the gain to changes in pump and signal power. The maximum gain of 20 dB is reached at a pump power of 330 nW. A plot of the amplified idler tone power at the spectrum analyzer versus input signal power is shown in Fig. 3.d. For this measurement, the KI-TWPA was operated as a frequency translating amplifier using the circuit shown in Fig. 5. The measurements were taken by varying the power of a signal tone at 550 MHz and measuring the idler tone power at 2.45 GHz on a spectrum analyzer. The pump tone was set to 1.5 GHz, and the measurements were taken for two different pump powers. The corresponding idler-to-idler gains measured for a pump power of -35 dBm and -35.6 dBm are 21 and 18 dB, respectively. Gain compression for a TWPA results from pump depletion and sets in when the output power of the signal and idler tones becomes large enough to significantly perturb the pump power.
@@ IV Noise Measurements
A Y-factor measurement was used to determine the noise added by the KI-TWPA and the total noise of the measurement system. A detailed diagram of the frequency translating mode noise measurement circuit is shown in fig. 5 in the appendix. Two 50 $\mathrm{\Omega}$ terminations were used as the noise sources, one at $T_{C}=10$ mK, thermalized to the mixing chamber (MXC) stage of a dilution refrigerator, and the other at the 3 K stage ($T_{H}=3.13$ K). A cryogenic relay switch at the input of the KI-TWPA allowed for switching between the ’hot’ and ’cold’ sources. A pair of diplexers at the input of the KI-TWPA were used to limit the noise power from the sources to the signal band of the KI-TWPA, inject the pump tone, and inject a separate test tone over the idler frequency range for measurement of the gain of the KI-TWPA. The idler frequency input had 40 dB attenuators at 10 mK and 4 K to ensure that the KI-TWPA was only presented with vacuum noise over the idler frequency range. A second pair of diplexers at the output of the KI-TWPA routed the signal band to a MXC stage termination and the pump tone to a separate termination on a higher temperature stage to prevent the MXC from heating. The output idler band tones were passed through an isolator, a cryogenic HEMT amplifier, and a room temperature amplifier and measured using a spectrum analyzer. The system noise in units of quanta is determined using the standard Y-factor relation
$N_{sys}=\frac{N_{\mathrm{H}}(\omega_{s})-YN_{\mathrm{C}}(\omega_{s})}{Y-1},$ (5)
where the $Y=P_{H}(\omega_{i})/P_{C}(\omega_{i})$ is the ratio of the power measured on the spectrum analyzer at the idler frequency with the hot and cold loads connected, and $N_{H,C}=\frac{1}{2}\coth(\hbar\omega_{s}/k_{b}T_{H,C})$. We also extract the contribution of the KI-TWPA to the measured noise using a modified Y-factor
@@
$A_{\mathrm{PA}}=\frac{N_{\mathrm{H}}(\omega_{s})-Y^{\prime}N_{\mathrm{C}}(\omega_{s})}{Y^{\prime}-1}.$ (7)
The derivation of this equation is discussed in the appendix. Fig. 4.a shows system noise and KI-TWPA added noise with the pump tone at 1.65 GHz. Fig. 4.b shows the KI-TWPA added noise for different pump frequencies, covering the 450 MHz –800 MHz signal band. The added noise reaches about 1 photon, slightly higher than the quantum limit of 1/2 photon, while the system noise reaches about 2 photons. The excess system noise is roughly consistent with the noise of the HEMT amplifier from the specifications ($T_{N}\approx 1.5$ K) and an estimate of the losses between the KI-TWPA and the HEMT.
@@ A.1 Device Fabrication
Appendix B Noise Measurement setup
@@ B.1 Noise Theory
$P_{\mathrm{C,H}}(\omega_{i})=B\,\hbar\omega_{s}\Big{[}N_{\mathrm{C,H}}(\omega_{s})+A^{\prime}_{PA}\Big{]}G_{si}G_{H}G_{R}$ (8)
$+\,B\,\hbar\omega_{i}\Big{[}N_{T}(\omega_{i})G_{ii}G_{H}G_{R}+A_{H}G_{H}G_{R}+A_{R}G_{R}\Big{]},$
where $B$ is the bandwidth, the $N_{\mathrm{C,H,T}}(\omega)=(1/2)\coth{(\hbar\omega/2kT_{\mathrm{C,H,T}})}$ are the noise quanta from the cold and hot loads and from the idler input termination expressed in photon units. $N_{T}(\omega_{i})$ is close to the vacuum state value of 0.5 photons. $G_{si}$ is the signal to idler power gain from eqn. 4, and $G_{H}$ and $G_{R}$ are the HEMT and room temperature amplifier gains. The $A_{H}$ and $A_{R}$ are the added noise of the HEMT and room temperature amplifiers in photon units. $A^{\prime}_{PA}$ is the added noise of the KI-TWPA in excess of the noise entering at the idler frequency, as the idler noise contribution is included separately in eqn. 8.
@@
$P_{\mathrm{off}}(\omega_{i})=B\,\hbar\omega_{i}\Big{[}\big{(}N_{T}(\omega_{i})+A_{H}\big{)}G_{H}G_{R}+A_{R}G_{R}\Big{]}.$ (9)
@@
$Y^{\prime}=\frac{N_{H}(\omega_{s})+A^{\prime}_{\mathrm{PA}}+N_{T}(\omega_{i})}{N_{C}(\omega_{s})+A^{\prime}_{\mathrm{PA}}+N_{T}(\omega_{i})}.$ (10)
@@
$A^{\prime}_{\mathrm{PA}}=\frac{N_{\mathrm{H}}(\omega_{s})-Y^{\prime}N_{\mathrm{C}}(\omega_{s})}{Y^{\prime}-1}-N_{\mathrm{T}}(\omega_{i}).$ (11)
@@
B.2 Frequency Translating Measurement Setup

### v2

@@ Abstract
We present the design and experimental characterization of a kinetic-inductance traveling-wave parametric amplifier (KI-TWPA) for sub-GHz frequencies. KI-TWPAs amplify signals through mixing processes supported by the nonlinear kinetic inductance of a superconducting transmission line. The device described here utilizes a compactly meandered TiN microstrip transmission line to achieve the length needed to amplify sub-GHz signals. It is operated in a frequency translating mode where the amplified signal tone is terminated at the output of the amplifier, and the idler tone at approximately 2.5 GHz is brought out of the cryostat. By varying the pump frequency, a gain of up to 22 dB was achieved in a tunable range from about 450 to 850 MHz. Use of TiN as the nonlinear element allows for a reduction of the required pump power by roughly an order of magnitude relative to NbTiN, which has been used for previous KI-TWPA implementations. This amplifier has the potential to enable high-sensitivity and high-speed measurements in a wide range of applications, such as quantum computing, astrophysics, and dark matter detection. 111© 2024. All rights reserved
@@ I Introduction
Kinetic inductance traveling-wave parametric amplifiers (KI-TWPAs)[1] with near quantum-limited noise performance and wideband gain have been developed recently and used for the readout of low-temperature detectors and qubits [2, 3, 4, 5]. The application of KI-TWPAs to direct dark matter searches is also gaining attention [6, 7]. The majority of these amplifiers have been designed for the 4-8 GHz band. However, a few technologies and experiments in quantum information science, astronomy, and fundamental physics can benefit from a near-quantum-noise-limited amplifier with gain at sub-GHz frequencies. For example, single-shot readout of spin-qubits operating below 1 GHz may be possible using a quantum-limited sub-GHz amplifier [8, 9, 10, 11]. Reducing the readout noise of nano-calorimeters and nano-bolometers would allow them to be sensitive to single photons in the microwave regime [12, 13, 14, 15, 16, 17]. Circuit quantum cavity electrodynamics with micro-mechanical hybrid systems can also benefit from sub-GHz quantum limited amplifiers [18, 19]. Far- and mid-IR MKID detectors operating in the single photon counting regime[21] could be improved with lower noise readout amplifiers operating around 1 GHz.
@@
This article presents results on a wide bandwidth sub-GHz KI-TWPA with sensitivity close to the quantum limit. The design of the amplifier draws on our earlier work[28, 29], but introduces two new innovations. The first departure from previous work is the use of a TiN microstrip line in place of NbTiN as the nonlinear transmission line, which significantly reduces the required pump power. Reduced pump power facilitates thermal control of the experimental setup at the low temperatures required to avoid thermal noise. For example, at 500 MHz, 1 photon corresponds to $hf/k_{\mathrm{B}}=24$ mK. KI-TWPA designs based on 4-wave mixing (4WM) typically needed a pump power of $\sim 10\,\mu$W. While the KI-TWPA itself dissipates only a fraction of the pump power, it is harder to control dissipation in the surrounding circuitry, including microwave isolator. The TiN device described here requires only $\sim 0.3\,\mu$W, which causes minimal heating at the base temperature of a dilution refrigerator.
@@ II Device Design
Similar to KI-TWPAs produced previously in our laboratory, the device reported here is based on a nonlinear transmission line (NTRL) structure with an inverted microstrip line geometry [28, 30, 29]. The microstrip conductor layer is patterned from a 50 nm thick TiN film. The dielectric of the microstrip is a 100 nm thick amorphous Si layer, and the sky plane is 200 nm thick niobium. The width of the microstrip line is 320 nm. Every 15 $\mu$m  a pair of microstrip open stubs of the same width and average length 120 $\mu$m  are connected to the central microstrip line. The NTRL is a slow-wave structure with propagation velocity $\sim 0.0064c$ and is tightly meandered, as shown in fig. 1a., across a 25 mm long chip. The 45 cm total length of the NTRL is several times longer than KI-TWPAs designed for > 4 GHz in order to roughly maintain the number of wavelengths ($\sim$300 pump wavelengths) for sufficient gain at the sub-GHz signal frequencies.
@@
Figure 2: a) Plot of the gain (pump on / pump off transmission) of the device with an applied -34.8 dBm pump tone at 1.6 GHz, measured using a VNA. The input and output frequencies are the same for this measurement. The dashed line is the gain calculated by integrating coupled mode equations.[28, 31] b) The device may also be operated as a frequency translating amplifier using a set of diplexers. At its input, the signal band is admitted while the device sees a termination over the idler band. At the output, the idler band is transmitted, and the signal band is terminated.
@@ III Performance
Figure 3: Measurements in subfigures a,b, and c were taken at 200 mK with a 20 dB attenuator at the 4 K stage. A 20 dB directional coupler was used at room temperature to couple the pump and signal tones into the cryostat. a) Measured signal band gain for a range of pump frequencies and all with a pump power of -34.8 dBm at the device. b) Corresponding idler band gain. c) Plot of the measured signal gain at 200 mK as a function of pump power for a pump tone at 1.6 GHz. The maximum gain near 20 dB corresponds to a pump power of 330 nW. At higher pump power the device switches to the normal state. d) A plot of the $\mathrm{P_{1dB}}$ and $\mathrm{IP_{3}}$ measurements is presented, taken with a pump tone at 1.6 GHz and a gain of $\sim$21.5 dB.
Applying a pumping tone below the stop band frequency results in parametric amplification via 4WM for weaker signal tones. The typical gain shown in fig. 2.a has three disjointed bands where gain is produced, similar to the device described in [29], but with a wider separation between bands. The low and high frequency bands reflect the frequencies at which the phase-matching criterion for 4WM is approximately satisfied:
@@
where $I_{p}$ is the pump current, the $k(\omega_{s,i,p})$ are the propagation constants at the signal, idler and pump frequencies, and $2\omega_{p}=\omega_{s}+\omega_{i}$. The central band reflects non-phase-matched gain around the pump frequency. The dashed line in the figure shows the gain calculated by integrating the coupled mode equations [28, 31] including third harmonic and upper sideband generation (tones at $3\omega_{p}$, $2\omega_{p}+\omega_{s}$ and $2\omega_{p}+\omega_{i}$). The film penetration depth in the model was adjusted to 625 nm to match the measured stop band frequency. To match the overall gain, the pump current was set to $I_{p}/I_{\ast}=0.117$, corresponding to 300 nW, which is close to the estimate of the experimental pump power.
@@
Fig. 3.c shows the response of the gain to changes in pump and signal power. The maximum gain of 20 dB is reached at a pump power of 330 nW. The 1 dB compression point ($\mathrm{P_{1dB}}$) was determined by gradually increasing the input power of a signal tone and measuring the corresponding output power using a spectrum analyzer. For this measurement, an input signal frequency of 675 MHz was selected, while a pump tone at 1.6 GHz with an input power of -35 dBm was applied to achieve a gain of $\sim$ 21.5 dB. The third-order intercept point ($\mathrm{IP_{3}}$) was measured using two signal tones at 675 MHz and 655 MHz, each with equal input power with the same pump conditions used in the $\mathrm{P_{1dB}}$ measurement. The input powers of the signal tones were gradually increased, and the corresponding output powers were recorded using a spectrum analyzer. Additionally, the output powers of the third-order intermodulation products ($f_{1}-f_{2}$) and ($f_{2}-f_{1}$) were measured to determine the third-order intercept point ($\mathrm{IP_{3}}$). At a gain of approximately 21.5 dB, the $\mathrm{P_{1dB}}$ was measured to be -71.1 dBm, and the $\mathrm{IP_{3}}$) was -59.4 dBm. A plot showing the compression point at different gain levels is included in Appendix C for further reference.
Gain compression for a 4WM TWPA results from pump depletion and sets in when the output power of the signal and idler tones becomes large enough to significantly perturb the pump power.
@@ IV Noise Measurements
A Y-factor measurement was used to determine the noise added by the KI-TWPA and the total noise of the measurement system. A detailed diagram of the frequency translating mode noise measurement circuit is shown in fig. 8 in the appendix. Two 50 $\mathrm{\Omega}$ terminations were used as the noise sources, one at $T_{\mathrm{cold}}=10$ mK, thermalized to the mixing chamber (MXC) stage of a dilution refrigerator, and the other at the 3 K stage ($T_{\mathrm{hot}}=3.13$ K). The hot termination is connected to the MXC stage through a superconducting cable with negligible loss over the signal band. A cryogenic relay switch at the input of the KI-TWPA allowed for switching between the ’hot’ and ’cold’ sources. A pair of diplexers at the input of the KI-TWPA were used to limit the noise power from the sources to the signal band of the KI-TWPA, inject the pump tone, and inject a separate test tone over the idler frequency range for measurement of the gain of the KI-TWPA. The idler frequency input had 40 dB attenuators at 10 mK and 3.13 K to ensure that the KI-TWPA was only presented with vacuum noise over the idler frequency range. A second pair of diplexers at the output of the KI-TWPA routed the signal band to a MXC stage termination and the pump tone to a separate termination on a higher temperature stage to prevent the MXC from heating. The output idler band tones were passed through an isolator, a cryogenic HEMT amplifier, and a room temperature amplifier and measured using a spectrum analyzer. The system noise in units of quanta is determined using the Y-factor relation
$N_{\mathrm{sys}}=\frac{N^{\prime}_{\mathrm{H}}(\omega_{s})-YN_{\mathrm{C}}(\omega_{s})}{Y-1},$ (5)
where the $Y=P_{\mathrm{H}}(\omega_{i})/P_{\mathrm{C}}(\omega_{i})$ is the ratio of the power measured on the spectrum analyzer at the idler frequency with the hot and cold loads connected and $N^{\prime}_{\mathrm{H}}(\omega_{s})=L_{1}N_{\mathrm{H}}(\omega_{s})+(1-L_{1})N_{\mathrm{T}}(\omega_{s})$ is the effective hot load noise seen by the TWPA through the input diplexer, which has transmission coefficient $L_{1}\approx 0.95$. $N_{\mathrm{H},\mathrm{C},\mathrm{T}}(\omega)=\frac{1}{2}\coth(\hbar\omega/2k_{b}T_{\mathrm{H},\mathrm{C},\mathrm{T}})$ are the thermal noise in photon units at the temperatures $T_{\mathrm{H}}$, $T_{\mathrm{C}}$ and the mixing chamber temperature $T_{\mathrm{T}}$. The contribution of the KI-TWPA to the measured noise can be extracted using a modified Y-factor
@@
$A_{\mathrm{PA}}=\frac{N^{\prime}_{\mathrm{H}}(\omega_{s})-Y^{\prime}N_{\mathrm{C}}(\omega_{s})}{Y^{\prime}-1}.$ (7)
The derivation of this equation is discussed in the appendix. Fig. 7.a shows system noise and KI-TWPA added noise with the pump tone at 1.65 GHz. Fig. 7.b shows the KI-TWPA added noise for different pump frequencies, covering the 450 MHz – 800 MHz signal band. The added noise reaches about 1 photon, slightly higher than the quantum limit of 1/2 photon, while the system noise reaches about 2 photons. The uncertainty in the hot load temperature is estimated to be ±$\pm 50$mK , leading to an uncertainty of ±$\pm 0.25$ photons in measured system noise. The excess system noise is roughly consistent with the noise of the HEMT amplifier from the specifications, $N_{\mathrm{HEMT}}\sim 50$ photons for $T_{N}\approx 1.5$ K, and an estimate of the losses between the KI-TWPA and the HEMT.
@@ Appendix A Design & Fabrication
[…]
Figure 5: The circuit diagram of the noise model, presented in Appendix C.1, illustrates the various noise contributions in the system. $N_{\mathrm{H}}(\omega_{s})$, $N_{\mathrm{C}}(\omega_{s})$, and $N_{\mathrm{T}}$ represent the thermal noise, expressed in photon units, corresponding to the temperatures $T_{\mathrm{H}}$, $T_{\mathrm{C}}$, and the mixing chamber temperature $T_{\mathrm{T}}$, respectively. The term $A^{\prime}_{PA}$ denotes the added noise of the TWPA, while $G_{si}$ and $G_{ii}$ represent the signal-to-idler and idler-to-idler gains, respectively. The loss between the noise sources and the device is characterized as $L_{1}(\omega_{s})$, while the loss between the device and the HEMT amplifier is denoted as $L_{2}(\omega_{i})$.
@@ A.1 Device Fabrication
Appendix B Dielectric Loss, Extraneous Mixing Products and Gain Compression
Fig 6.a shows the measured insertion loss of the housed KI-TWPA device at a temperature of 200 mK. The loss in the signal band ranges from 0.7 to 1 dB, whereas the loss in the idler band is approximately 1.5 dB. As single-layer resonators made from identically prepared TiN films have $Q$ factors in excess of $10^{6}$, it is likely that the majority of the loss is from the aSi dielectric layer, with possibly a small contribution from the normal metal transitions structures between the microwave connectors and the chip.
Fig 6.b is the measured output spectrum of the pumped device at 200 mK for a low power input signal and pump frequency of 1.6 GHz. Higher-frequency components, including harmonics such as $3\omega_{p}$, $2\omega_{p}$, and other nonlinear mixing products, are clearly identifiable in the output spectrum of the device. The amplitudes of the third harmonic and “upper sideband” tones at $4\omega_{p}-\omega_{s,i}$ (only $4\omega_{p}-\omega_{i}$ lies in the frequency range shown) agree well with the prediction of the coupled mode equations[28] and the calculated dispersion. The small amplitude of the third harmonic relative to the pump and of the upper sidebands relative to the signal, indicate that the dispersion engineering of the device is working as intended. A stronger coupling to the upper sideband frequencies could increase the noise by mixing in vacuum fluctuations from those frequencies. The level of the second harmonic is higher than is expected from the calculation, and the origins of the unlabeled tones are unknown.
Figure 6: a) Plot of measured insertion loss of the device plus the package as a function of frequency. b) Measured output power of the device at low signal power.
The compression point of the amplifier was also measured for an input signal frequency of 675 MHz. The pump tone was applied at 1.6GHz, and the pump power was adjusted to attain gain levels of 10 dB, 16.5 dB, and 21.5 dB. The output power of the signal tone was measured as a function of the signal input power level.
Figure 7: Plot of the measured output power at the device of a test tone as a function of the test tone’s input power. The measurements were performed with a pump tone at 1.6 GHz and signal gains of 10 dB, 16.5 dB, and 21.5 dB.
Appendix C Noise Measurement setup
@@ C.1 Noise Theory
$P_{\mathrm{C},\mathrm{H}}$ $(\omega_{i})=$ (8)
[($\bigl{[}\bigl{(}\{\hbar\omega_{s}G_{si}\left[N^{\prime}_{\mathrm{C},\mathrm{H}}(\omega_{s})+A^{\prime}_{PA}\right]+\hbar\omega_{i}G_{ii}N_{\mathrm{T}}(\omega_{i})\}L_{2}$
$+\hbar\omega_{i}(\{1-L_{2}\}N_{\mathrm{T}}(\omega_{i})+A_{\mathrm{HEMT}})\bigr{)}G_{\mathrm{HEMT}}+A_{R}\bigr{]}G_{R}B$
where $B$ is the bandwidth, $1-L_{2}\approx 0.3$ is the loss between the KI-TWPA and the HEMT amplifier and $1-L_{1}\approx 0.05$ is the loss between the KI-TWPA and the hot termination. The small losses corresponding to $L_{1}$ and $L_{2}$ have little effect on analysis, but are included for completeness. At 600 MHz and the operating temperature of 0.012 K, $N_{T}(\omega_{i})=0.5$ and $N_{T}(\omega_{s})=0.6$, close to the vacuum level. $G_{si}$ is the signal to idler power gain from eqn. 4, and $G_{\mathrm{HEMT}}$ and $G_{\mathrm{R}}$ are the HEMT and room temperature amplifier gains. The $A_{\mathrm{HEMT}}$ and $A_{\mathrm{R}}$ are the added noise of the HEMT and room temperature amplifiers in photon units. $A^{\prime}_{\mathrm{PA}}$ is the added noise of the KI-TWPA in excess of the noise entering at the idler frequency, as the idler noise contribution is included separately in eqn. C.1. $N^{\prime}_{\mathrm{H}}=N_{\mathrm{H}}L_{1}+N_{\mathrm{T}}(\omega_{s})(1-L_{1})$ is the noise from the hot load corrected for the loss $1-L_{1}$, and $N^{\prime}_{\mathrm{C}}=N_{\mathrm{C}}$ because the temperatures of the cold load and connecting lossy elements are the same. The contributions to the total output noise are diagrammed in fig. 5.
@@
$P_{\mathrm{off}}(\omega_{i})=\hbar\omega_{i}\Big{[}\big{(}N_{\mathrm{T}}(\omega_{i})+A_{\mathrm{HEMT}}\big{)}G_{\mathrm{HEMT}}+A_{\mathrm{R}}\Big{]}G_{\mathrm{R}}\,B.$ (9)
@@
$Y^{\prime}=\frac{N^{\prime}_{\mathrm{H}}(\omega_{s})+A^{\prime}_{\mathrm{PA}}+N_{\mathrm{T}}(\omega_{i})}{N^{\prime}_{\mathrm{C}}(\omega_{s})+A^{\prime}_{\mathrm{PA}}+N_{\mathrm{T}}(\omega_{i})}.$ (10)
@@
$A^{\prime}_{\mathrm{PA}}=\frac{N^{\prime}_{\mathrm{H}}(\omega_{s})-Y^{\prime}N^{\prime}_{\mathrm{C}}(\omega_{s})}{Y^{\prime}-1}-N_{\mathrm{T}}(\omega_{i}).$ (11)
@@
C.2 Frequency Translating Measurement Setup